# A Nonparametric Discrete Hawkes Model with a Collapsed Gaussian-Process Prior

### v1

@@ 1 Introduction
Hawkes processes are self-exciting stochastic models in which the probability of an event occurring increases in response to past occurrences. Originally introduced by (hawkes1971, ), they have been widely adopted across disciplines such as seismology (ogata1988, ), finance (bacry2015, ), criminology (mohler2011, ), social networks (zhou2013, ; du2016, ), and epidemiology (reinhart2018, ; browning2021, ). Their appeal lies in the ability to model bursty, temporally clustered behavior via an additive decomposition of the event rate into baseline and excitation components.
Much of the literature has focused on continuous-time Hawkes processes, where the conditional intensity evolves over real-valued time and is typically expressed as the sum of a deterministic baseline and a parametric excitation kernel. Common excitation functions include exponential and power-law decay due to their simplicity and interpretability (ogata1988, ). More recent work proposes flexible alternatives using histogram-based kernels (lewis2011, ), basis expansions (zhou2013, ), neural network parameterizations (mei2017, ), and Gaussian process priors (adams2009, ; lloyd2015, ), with fully Bayesian approaches for latent network inference (linderman2014, ). In particular, Gaussian process based intensities have been developed in continuous time (e.g. (malemshinitski2021, )).
@@
To address these limitations, (browning2022, ) introduced a nonparametric DHP using a random histogram prior over the excitation kernel with trans-dimensional MCMC for inference. While this approach allows data-driven excitation structure, the use of a fixed intercept for the baseline may limit its ability to capture smooth or evolving background dynamics. Moreover, the sampling scheme can be costly for long processes.
@@ 2.1 Discrete Hawkes Process
The discrete-time Hawkes process (DHP) is the discrete analogue of the continuous-time Hawkes process originally introduced by (hawkes1971, ). In continuous time, the intensity function defines the instantaneous rate of event occurrence and typically increases immediately following a self-exciting or clustering phenomenon. In contrast, the DHP operates over discrete time steps, $t\in\mathbb{N}$, and models the expected number of events per interval using a combination of exogenous and endogenous contributions.
@@ 2.1.1 Branching Representation of the Discrete Hawkes Process
The discrete Hawkes model above is often interpreted through the Hawkes cluster representation: events arise either as immigrants from the baseline $\mu(t)$ or as offspring triggered by past events, with integer lags distributed according to $\Phi(d)$ (hawkes1971, ; hawkes1974, ). In this view, bursts form around immigrant arrivals as offspring generate further descendants. We call this the branching process interpretation.
@@ 2.2 Proposed Model
Standard inference in Hawkes processes typically involves estimating the baseline and excitation functions either through maximum likelihood using parametric forms, or via sampling-based Bayesian approaches. In the discrete-time setting, parametric models often assume constant or periodic baselines and simple excitation forms (e.g., exponential decay (browning2021, )), which limit flexibility. Fully Bayesian approaches with nonparametric priors exist but tend to be computationally intensive, especially when jointly estimating multiple latent functions. In contrast, our approach collapses the GP priors over baseline and excitation into a single prior over the latent intensity trajectory, enabling efficient MAP inference with scalable optimization. This not only improves interpretability but also avoids expensive sampling or marginalization steps, making the method suitable for large-scale temporal data. To this end, we propose a nonparametric extension of the DHP that models both components using Gaussian process (GP) priors. While several works have explored GP-based intensity modeling in the continuous-time Hawkes framework, see for example adams2009 ; lloyd2015 ; malemshinitski2021 , to our knowledge, there are currently no analogous GP-based formulations for discrete-time Hawkes processes. Our work addresses this gap by developing a discrete-time, nonparametric model that retains the interpretability of parametric counterparts. We first introduce the Gaussian Process.
@@ 2.2.1 Gaussian processes
In what follows, we place GP priors on both the baseline $b(t)$ and the excitation kernel $f(d)$ of the discrete Hawkes model. This allows interpretable, data-driven intensity dynamics without committing to fixed parametric forms. For a comprehensive introduction, see (rasmussen2006gaussian, ).
@@ Discussion.
The combination of (3) and (4) provides the necessary inductive bias for identifiability: repeated seasonal structure and trend are captured by $b(t)$, whereas $f(d)$ explains short- to medium-range self-excitation with a principled decay built into the prior malemshinitski2021 (i.e., without an explicit Gaussian envelope factor $e^{-\alpha d}$ multiplied the excitation function).
@@ 3 Inference
Our inference procedure centers on the latent intensity function $l(t)$, defined in Equation (2), which we treat as the primary object of estimation. Since both the baseline $b(t)$ and the excitation kernel $f(d)$ are endowed with Gaussian process priors, the resulting $l(t)$ inherits a well-defined Gaussian process prior induced by these components (rasmussen2006gaussian, ). Rather than estimating $b(t)$ and $f(d)$ jointly, we integrate them out analytically to form a collapsed prior over $l(t)$. This yields a single GP prior over the entire latent trajectory, improving computational efficiency while preserving the underlying structure.
@@ 3.1 Collapsed GP Prior
This formulation avoids the need to represent $b(t)$ and $f(d)$ explicitly during inference. This is analogous to GP-modulated Poisson processes in continuous time malemshinitski2021 , adapted here to the discrete-time Hawkes setting with a collapsed prior. Unlike malemshinitski2021 , which imposes an explicit exponential envelope together with a stationary RBF kernel for the excitation (i.e., contributions of the form $e^{-\alpha d}$ and covariance of the GP prior of the excitation function $k_{f}(d,d^{\prime})=\sigma_{f}^{2}\exp\!\{-(d-d^{\prime})^{2}/(2\ell_{f}^{2})\}$), we remove the deterministic envelope from the mean structure and encode attenuation and smoothness inside the GP prior for $f$.
@@ 3.4 Computational Complexity and Efficiency
    where $K_{U}$ is the stationary RBF kernel on the uniform grid. Because the grid is uniform, $K_{U}$ is (block-)Toeplitz/circulant and supports $\mathcal{O}(M\log M)$ MVMs via FFT (see xia2007toeplitz ). Since $W$ is very sparse, $W\boldsymbol{v}$ and $W^{\top}\boldsymbol{v}$ cost $\mathcal{O}(T)$.
@@ 4.2 Real Data
We evaluate GP-DHP and four benchmark parametric discrete-time Hawkes models on two real-world count datasets. For the benchmarks, we fix the excitation function to a negative-binomial form (a commonly used parametric form of the excitation function (white2012terrorism, )), whereas GP-DHP models the excitation nonparametrically via a Gaussian process prior. All benchmark models share the same negative-binomial excitation kernel:
@@
which has been used in recent infectious disease applications, including modeling COVID-19 mortality browning2021 . The more flexible negative binomial kernel has also been used in terrorism modeling white2012terrorism .
@@
We analyze a 21-year time series of daily terrorist incidents within the United States, drawn from the Global Terrorism Database (GTD) gtd . The data spans from 1972 onward. All events are aggregated to daily resolution and partitioned into training (1972–1981), validation (1982–1987), and test (1988–1992) intervals.
@@
Self-exciting models offer a principled alternative. Prior work by (white2012terrorism, ) applied a discrete Hawkes process with a negative binomial excitation kernel to terrorism data in Southeast Asia, demonstrating how such models can recover meaningful structure in burst-prone time series and yield interpretable quantities such as volatility and resilience.
@@ 4.2.2 Cryptosporidiosis
We analyze a 365-week series of weekly Cryptosporidiosis case counts, available via the tscount package in R tscount . Cryptosporidiosis is a gastrointestinal disease caused by the protozoan parasite Cryptosporidium, which is typically spread through ingestion of contaminated water. While low-level background transmission persists in many regions, the disease is most notable for its sporadic and localized outbreaks. One of the most severe documented events occurred in Milwaukee in 1993, where over 400,000 residents were affected due to a contaminated municipal water supply hunter2003epidemiology ; yoder2010cryptosporidiosis . These outbreak dynamics are characterized by sharp, short-term increases in reported cases, followed by a rapid return to low endemic levels.
Such temporal patterns are challenging for classical autoregressive count models, such as Poisson or INGARCH-type models fokianos2009poisson ; fokianos2011log , which assume dependence based on past counts but do not directly model the self-exciting nature of outbreaks. These models often struggle to distinguish between sustained background intensity and transient, event-driven clustering.
@@
Recent work by (browning2021, ) demonstrated the use of discrete Hawkes models with geometric excitation kernels for modeling daily COVID-19 mortality in multiple countries. The geometric kernel is a special case of the negative binomial excitation used here (specifically, when $r=1$), and has proven effective for modeling short-memory self-excitation in epidemic settings. By generalizing to a full negative binomial kernel, GP-DHP can flexibly adapt to varying outbreak shapes and durations. On top of this, the branching ratio for the excitation function was $\sum_{d\geq 1}\max\{\Phi(d),0\}=0.86$. Since the branching ratio is less than one, we get a stable process.

### v2

@@ 1 Introduction
Hawkes processes are self-exciting stochastic models in which the probability of an event occurring increases in response to past occurrences. Originally introduced by (Hawkes, 1971), they have been widely adopted across disciplines such as seismology (Ogata, 1988), finance (Bacry et al., 2015), criminology (Mohler et al., 2011), social networks (Zhou et al., 2013; Du et al., 2016), and epidemiology (Reinhart, 2018; Browning et al., 2021). Their appeal lies in the ability to model bursty, temporally clustered behavior via an additive decomposition of the event rate into baseline and excitation components.
Much of the literature has focused on continuous-time Hawkes processes, where the conditional intensity evolves over real-valued time and is typically expressed as the sum of a deterministic baseline and a parametric excitation kernel. Common excitation functions include exponential and power-law decay due to their simplicity and interpretability (Ogata, 1988). More recent work proposes flexible alternatives using histogram-based kernels (Lewis and Mohler, 2011), basis expansions (Zhou et al., 2013), neural network parameterizations (Mei and Eisner, 2017), and Gaussian process priors (Adams et al., 2009; Lloyd et al., 2015), with fully Bayesian approaches for latent network inference (Linderman and Adams, 2014). In particular, Gaussian process based intensities have been developed in continuous time (e.g. (Malem-Shinitski et al., 2022)).
@@
To address these limitations, (Browning et al., 2022) introduced a nonparametric DHP using a random histogram prior over the excitation kernel with trans-dimensional MCMC for inference. While this approach allows data-driven excitation structure, the use of a fixed intercept for the baseline may limit its ability to capture smooth or evolving background dynamics. Moreover, the sampling scheme can be costly for long processes.
@@ 2.1 Discrete Hawkes Process
The discrete-time Hawkes process (DHP) is the discrete analogue of the continuous-time Hawkes process originally introduced by (Hawkes, 1971). In continuous time, the intensity function defines the instantaneous rate of event occurrence and typically increases immediately following a self-exciting or clustering phenomenon. In contrast, the DHP operates over discrete time steps, $t\in\mathbb{N}$, and models the expected number of events per interval using a combination of exogenous and endogenous contributions.
@@ 2.1.1 Branching Representation of the Discrete Hawkes Process
The discrete Hawkes model above is often interpreted through the Hawkes cluster representation: events arise either as immigrants from the baseline $\mu(t)$ or as offspring triggered by past events, with integer lags distributed according to $\Phi(d)$ (Hawkes, 1971; Hawkes and Oakes, 1974). In this view, bursts form around immigrant arrivals as offspring generate further descendants. We call this the branching process interpretation.
@@ 2.2 Proposed Model
Standard inference in Hawkes processes typically involves estimating the baseline and excitation functions either through maximum likelihood using parametric forms, or via sampling-based Bayesian approaches. In the discrete-time setting, parametric models often assume constant or periodic baselines and simple excitation forms (e.g., exponential decay (Browning et al., 2021)), which limit flexibility. Fully Bayesian approaches with nonparametric priors exist but tend to be computationally intensive, especially when jointly estimating multiple latent functions. In contrast, our approach collapses the GP priors over baseline and excitation into a single prior over the latent intensity trajectory, enabling efficient MAP inference with scalable optimization. This not only improves interpretability but also avoids expensive sampling or marginalization steps, making the method suitable for large-scale temporal data. To this end, we propose a nonparametric extension of the DHP that models both components using Gaussian process (GP) priors. While several works have explored GP-based intensity modeling in the continuous-time Hawkes framework, see for example Adams et al. (2009); Lloyd et al. (2015); Malem-Shinitski et al. (2022), to our knowledge, there are currently no analogous GP-based formulations for discrete-time Hawkes processes. Our work addresses this gap by developing a discrete-time, nonparametric model that retains the interpretability of parametric counterparts. We first introduce the Gaussian Process.
@@ 2.2.1 Gaussian processes
In what follows, we place GP priors on both the baseline $b(t)$ and the excitation kernel $f(d)$ of the discrete Hawkes model. This allows interpretable, data-driven intensity dynamics without committing to fixed parametric forms. For a comprehensive introduction, see (Rasmussen and Williams, 2006).
@@ Discussion.
The combination of (3) and (4) provides the necessary inductive bias for identifiability: repeated seasonal structure and trend are captured by $b(t)$, whereas $f(d)$ explains short- to medium-range self-excitation with a principled decay built into the prior Malem-Shinitski et al. (2022) (i.e., without an explicit Gaussian envelope factor $e^{-\alpha d}$ multiplied the excitation function).
@@ 3 Inference
Our inference procedure centers on the latent intensity function $l(t)$, defined in Equation (2), which we treat as the primary object of estimation. Since both the baseline $b(t)$ and the excitation kernel $f(d)$ are endowed with Gaussian process priors, the resulting $l(t)$ inherits a well-defined Gaussian process prior induced by these components (Rasmussen and Williams, 2006). Rather than estimating $b(t)$ and $f(d)$ jointly, we integrate them out analytically to form a collapsed prior over $l(t)$. This yields a single GP prior over the entire latent trajectory, improving computational efficiency while preserving the underlying structure.
@@ 3.1 Collapsed GP Prior
This formulation avoids the need to represent $b(t)$ and $f(d)$ explicitly during inference. This is analogous to GP-modulated Poisson processes in continuous time Malem-Shinitski et al. (2022), adapted here to the discrete-time Hawkes setting with a collapsed prior. Unlike Malem-Shinitski et al. (2022), which imposes an explicit exponential envelope together with a stationary RBF kernel for the excitation (i.e., contributions of the form $e^{-\alpha d}$ and covariance of the GP prior of the excitation function $k_{f}(d,d^{\prime})=\sigma_{f}^{2}\exp\!\{-(d-d^{\prime})^{2}/(2\ell_{f}^{2})\}$), we remove the deterministic envelope from the mean structure and encode attenuation and smoothness inside the GP prior for $f$.
@@ 3.4 Computational Complexity and Efficiency
    where $K_{U}$ is the stationary RBF kernel on the uniform grid. Because the grid is uniform, $K_{U}$ is (block-)Toeplitz/circulant and supports $\mathcal{O}(M\log M)$ MVMs via FFT (see Xia (2007)). Since $W$ is very sparse, $W\boldsymbol{v}$ and $W^{\top}\boldsymbol{v}$ cost $\mathcal{O}(T)$.
@@ 4.2 Real Data
We evaluate GP-DHP and four benchmark parametric discrete-time Hawkes models on two real-world count datasets. For the benchmarks, we fix the excitation function to a negative-binomial form (a commonly used parametric form of the excitation function (Porter and White, 2012)), whereas GP-DHP models the excitation nonparametrically via a Gaussian process prior. All benchmark models share the same negative-binomial excitation kernel:
@@
which has been used in recent infectious disease applications, including modeling COVID-19 mortality Browning et al. (2021). The more flexible negative binomial kernel has also been used in terrorism modeling Porter and White (2012).
@@
We analyze a 21-year time series of daily terrorist incidents within the United States, drawn from the Global Terrorism Database (GTD) for the Study of Terrorism and to Terrorism) . The data spans from 1972 onward. All events are aggregated to daily resolution and partitioned into training (1972–1981), validation (1982–1987), and test (1988–1992) intervals.
@@
Self-exciting models offer a principled alternative. Prior work by (Porter and White, 2012) applied a discrete Hawkes process with a negative binomial excitation kernel to terrorism data in Southeast Asia, demonstrating how such models can recover meaningful structure in burst-prone time series and yield interpretable quantities such as volatility and resilience.
@@ 4.2.2 Cryptosporidiosis
We analyze a 365-week series of weekly Cryptosporidiosis case counts, available via the tscount package in R Liboschik et al. (2017). Cryptosporidiosis is a gastrointestinal disease caused by the protozoan parasite Cryptosporidium, which is typically spread through ingestion of contaminated water. While low-level background transmission persists in many regions, the disease is most notable for its sporadic and localized outbreaks. One of the most severe documented events occurred in Milwaukee in 1993, where over 400,000 residents were affected due to a contaminated municipal water supply Hunter et al. (2003); Yoder et al. (2010). These outbreak dynamics are characterized by sharp, short-term increases in reported cases, followed by a rapid return to low endemic levels.
Such temporal patterns are challenging for classical autoregressive count models, such as Poisson or INGARCH-type models Fokianos et al. (2009); Fokianos and Tjøstheim (2011), which assume dependence based on past counts but do not directly model the self-exciting nature of outbreaks. These models often struggle to distinguish between sustained background intensity and transient, event-driven clustering.
@@
Recent work by (Browning et al., 2021) demonstrated the use of discrete Hawkes models with geometric excitation kernels for modeling daily COVID-19 mortality in multiple countries. The geometric kernel is a special case of the negative binomial excitation used here (specifically, when $r=1$), and has proven effective for modeling short-memory self-excitation in epidemic settings. By generalizing to a full negative binomial kernel, GP-DHP can flexibly adapt to varying outbreak shapes and durations. On top of this, the branching ratio for the excitation function was $\sum_{d\geq 1}\max\{\Phi(d),0\}=0.86$. Since the branching ratio is less than one, we get a stable process.